Fine the impedance of a series RLC circuit if the inductive reactance, capacitive reactance and resistance are 184 ohms, 144 ohms and 30 ohms respectively. Also calculate the phase angle between voltage and current
1
Expert's answer
2020-02-24T10:20:54-0500

When connected in series, the component impedances are added together

(1) "Z=R+iw\cdot L+ \frac{1}{iw \cdot C}"

To determine the phase angle between voltage and current, impedance should be represented in the Euler's form

"Z=|Z|\cdot e^{iarg(Z)}"

Determine first real and imaginary parts of the impedance (1) "Z=R+iX" . We can see reactance"X=Lw-\frac{1}{Cw}" . Thus

"|Z|=\sqrt{R^2+X^2}=\sqrt{R^2+(Lw-1/Cw)^2}=\sqrt{30^2+(184-144)^2}=50\Omega"


"arg(Z)=arctan(\frac{X}{R})=arctan(\frac{Lw-\frac{1}{Cw}}{R})=arctan(\frac{40}{30})=53\degree"


Answer: The phase angle between voltage and current "{\displaystyle \theta } =53\degree"